# AtlasD: Automatic Local Symmetry Discovery

### v1

@@ Abstract
Existing symmetry discovery methods predominantly focus on global transformations across the entire system or space, but they fail to consider the symmetries in local neighborhoods. This may result in the reported symmetry group being a misrepresentation of the true symmetry. In this paper, we formalize the notion of local symmetry as atlas equivariance. Our proposed pipeline, automatic local symmetry discovery (AtlasD), recovers the local symmetries of a function by training local predictor networks and then learning a Lie group basis to which the predictors are equivariant. We demonstrate AtlasD is capable of discovering local symmetry groups with multiple connected components in top-quark tagging and partial differential equation experiments. The discovered local symmetry is shown to be a useful inductive bias that improves the performance of downstream tasks in climate segmentation and vision tasks.
@@ 2 Related Work
Equivariant Neural Networks. Equivariant neural networks use known symmetry as an inductive bias when fitting a model. Group equivariant CNN extends the translational equivariance of a CNN to rotations and reflection using group theory (Cohen & Welling, 2016). Other works focus on designing networks that are equivariant to a wider range of transformations; particularly, $\mathrm{E}(2)$ transformations on Euclidean plane (Weiler & Cesa, 2019), rotations on the sphere (Cohen et al., 2018), Lorentz group transformations (Gong et al., 2022), and $\mathrm{E}(n)$ transformations in higher-dimensional spaces (Satorras et al., 2021). These works focus on global transformations and require prior knowledge of the symmetry. In contrast, our pipeline focuses on discovering local symmetry.
@@
Symmetry discovery works also consider slightly varied problem settings. Some authors seek to find a subset of possible symmetries (Benton et al., 2020; Romero & Lohit, 2022). Others consider the case where the group acts on the latent space instead of the feature space (Yang et al., 2024; Gabel et al., 2023; Keurti et al., 2023; Koyama et al., 2023). Still, these papers mainly focus on global transformations. More relevant to local symmetry discovery is the work by Decelle et al. (2019), which attempts to see if two datapoints are related by a particular local transformation. This is distinct from AtlasD, where we characterize the local symmetry group in an interpretable manner.
Comparison with State-of-the-Art. We briefly compare AtlasD with existing state-of-the-art models. LieGAN, LieGG, and AtlasD are all capable of discovering continuous equivariances, but LieGG requires significant memory and LieGAN uses adversarial methods, which can be complex to train. Moreover, while LieGG cannot discover discrete symmetries entirely and LieGAN can only discover those with positive determinants, AtlasD proves capable of discovering orientation-reversing discrete symmetries. The largest difference, however, is the implicit hypothesis space of each model. The hypothesis space is a crucial component of a discovery method as it determines the set of symmetries that can be recovered. LieGG and LieGAN only consider global symmetries, but AtlasD can also discover local ones.
@@ 4 AtlasD: Automatic Local Symmetry Discovery
Existing discovery methods are fundamentally limited in that they completely ignore local symmetries. To address this problem, we first formulate atlas equivariance as a definition of local symmetry. Then, we detail our methodology for discovering local symmetry in the form of a Lie algebra basis and component group. Finally, we highlight theoretical connections to existing work as well as implementation notes.
@@ 4.1 Atlas Equivariance
To provide an intuition of local symmetry, we highlight the heat equation on a torus in Figure 2. Consider the time-stepping function that evolves the current state of the system for some fixed time interval. If we focus only on a neighborhood of the input feature field and its corresponding neighborhood in the output field, a local rotation in the input results in an identical local rotation to the output.
To define local symmetry formally, assume we have a map $\Phi$ that transforms an input feature field $F_{\text{in}}:{\mathcal{M}}\to\mathbb{R}^{\text{d}_{\text{in}}}$ to an output field $F_{\text{out}}:{\mathcal{M}}\to\mathbb{R}^{\text{d}_{\text{out}}}$. Then, suppose $\mathcal{A}$ is an atlas on ${\mathcal{M}}$ given by a finite collection of charts $\{(U_{c},\varphi_{c})\}_{c=1}^{N}$. We can relocate a feature field on ${\mathcal{M}}$ restricted to the neighborhood $U_{c}$ to a flat Euclidean space by pulling back over $\varphi_{c}^{-1}$:
@@
We define local (atlas) equivariance for task functions where the output signal depends locally on the input signal. We formalize this as $\mathcal{A}$ locality. The intuitive notion of a local function is that, under an appropriate choice of atlas, we can fully reconstruct the output field in any individual chart $U_{c}$ solely from the input field along the same chart. Thus, we say $\Phi$ is $\mathcal{A}$ atlas local if we are able to decompose it into various $\Phi_{c}$, where $\Phi_{c}$ is a map between the pullback of the $c$th chart of the input and output feature fields. We sometimes refer to $\{\Phi_{c}\}$ as the localized functions of $\Phi$. Formally,
@@
Here, the restriction of the feature field to the subset $\varphi_{c}(U_{c})$ indicates that we are not particular about the output of $\Phi_{c}$ outside the projected chart. For these atlas local functions, it is possible to consider the symmetry transformations that operate within local neighborhoods. We formalize this notion of local symmetry as follows.
@@ 4.2 Atlas Equivariance Discovery
In our problem setup, we assume we have a dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$ that models some unknown function $\Phi:\mathcal{X}\to\mathcal{Y}$ between feature fields on the same manifold. We further assume that a suitable atlas $\mathcal{A}$ for the problem is known. We aim to find the maximal matrix Lie group that $\Phi$ is $\mathcal{A}$ atlas equivariant to.
We propose AtlasD, automatic local symmetry discovery, to tackle the problem. First, we must model individual localized functions with neural networks or other differentiable oracles. The exact method is unique to each task, but is generally a simple regression problem. Further details are available in Appendix C. In contrast to global discovery techniques, we emphasize that the individual neural networks are localized maps rather than functions over the entire manifold.
We then find the equivariance group of the localized functions. There are important differences between our procedure and global symmetry discovery. In our setting, a group element must only act on a local chart, rather than the full space. Likewise, a core tenet of local symmetry is that we can apply different transformations to different regions of the feature field. Care must be taken, therefore, that the actions on different charts are truly independent.
@@
  Train a predictor network $\Phi_{c}$ for each chart $(U_{c},\varphi_{c})$ under the loss $\mathcal{L}\left(\Phi_{c}((\varphi_{c}^{-1})^{*}X_{i}),(\varphi_{c}^{-1})^{*}Y_{i}\right)$
  Given the trained predictors $\{\Phi_{c}\}$, discover the Lie algebra basis $\{B_{i}\}$Section 4.2.1}
  Given the trained predictors $\{\Phi_{c}\}$ and the discovered Lie algebra basis $\{B_{i}\}$, find the coset representatives $\{C_{\ell}\}$Section 4.2.2}
@@ 4.3 Connection to Gauge Equivariant CNN
A related notion of local symmetry is introduced by Cohen et al. (2019) when defining gauge equivariant CNNs. In short, gauge equivariance implies that one should be able to arbitrarily orient the local coordinate systems used to define features and compute convolutions. Hence, it is a property of the network modeling the task function, rather than a property of the task function itself. This is a notable difference compared to our work, where the atlas equivariance group is intrinsic to the system and can be discovered.
@@ 5.1 Global Symmetry Comparison
The goal is to classify between top quark and lighter quarks jets present in the Top Quark Tagging Reference Dataset (Kasieczka et al., 2019). The dataset contains 2M observations, consisting of four-momentum of up to $200$ particle jets. The classification is invariant to the entire Lorentz group $\mathrm{O}(1,3)$, which we will try to discover. We use our infinitesimal generator discovery pipeline to learn the invariances of the predictor. We seed our basis with $7$ generators. In Figure 4, we show that the discovered basis matches closely with that of $\mathrm{SO}^{+}(1,3)$, the identity component of the Lorentz group. Moreover, computing the invariant tensor using the method from Yang et al. (2023), we find that the invariant tensor has a cosine correlation of $0.9996$ with the ground truth Minkowski tensor $\text{diag}(-1,1,1,1)$. This is a strong result that is slightly superior to LieGAN’s cosine correlation of $0.9975$.
Figure 4: AtlasD discovers $\mathrm{O}^{+}(1,3)$ in the top tagging task. Here, each red and blue heatmap denotes a Lie algebra basis element. For each generator, the value of its entries are depicted by the individual colors. Generators 0, 1, 2 correspond to $\mathrm{SO}(3)$ rotation and generators 4, 5, 6 indicate boosts. The pink and green heatmap displays the computed invariant metric. The final two heatmaps depict the discovered representatives of the identity and parity components.
We then try to discover the various cosets of the symmetry group. We seed our discovery process with $K=256$ matrices. In the dataset, the time component of all momenta are positive, and hence it is difficult to find the time-reversal generator of the Lorentz group. In the last two heatmaps of Figure 4, we discover a parity transformation, which means that the entire learned symmetry group is $\mathrm{O}^{+}(1,3)$. In contrast, LieGAN cannot discover orientation-reversing transformations and hence only reports the identity component $\mathrm{SO}^{+}(1,3)$.
@@ 5.2 Partial Differential Equation
(a) Example input and output field for heat experiment.
(b) Depiction of atlases for heat experiment. Each blue or red square represents an individual chart.
Figure 5: Illustration of PDE experiment.
@@
In Section D.4, we compare results with LieGG.
Figure 6: PDE discovered symmetry. The results for the first and second atlas are depicted in the top and bottom row, respectively. In each case, the leftmost entry is the discovered infinitesimal generator, and the right two columns are the discovered cosets.
@@ 5.3 MNIST on Sphere
To highlight the benefits of using our learned results in downstream models, we design a projected MNIST segmentation task. In this experiment, we project three digits from the MNIST dataset onto a sphere (Figure 7). Before projection, each image is randomly rotated up to $60$ degrees clockwise or counterclockwise. The goal of the model is to classify each pixel as either the background or its numeric value. Although the rotation of the digits adds a local symmetry, there is no continuous global symmetry since the position of each of the three digits is fixed. A natural atlas to use in this problem is assigning a single chart to the region of each digit. We then train a predictor for each of the three charts using CNNs. In the discovery process, we seed our model with a single infinitesimal generator. To demonstrate the benefit of considering local symmetry, we compare our results against a modified LieGAN that represents global symmetries as subgroups of $\mathrm{SO}(3)$.
Figure 7: MNIST experiment setup. (a) The input feature fields (top) are given by three digits rotated then projected onto the equator of a sphere. To construct the output feature fields (bottom), the model must label each pixel as either background or its numeric value if it is a part of a number. (b) We highlight two of the charts used in our atlas.
@@ 5.4 ClimateNet
We use an atlas that has $4$ charts spread through the surface of the earth. When we seeded our model with $1,2,$ or $3$ infinitesimal generators, we find that the resultant basis is not similar across consecutive runs. This suggests that the symmetry group is actually $4$-dimensional. To confirm this, we plot a chart predictor’s output after applying various linear actions in Figure 10. The figure highlights that the predictor is mildly equivariant to a wide range of actions. In fact, Figure 9 demonstrates comparable magnitudes of the $4$ generators. All of these points to evidence that the atlas equivariance group is $\mathrm{GL}^{+}(2)$.
@@
Figure 9: AtlasD discovers $\mathrm{GL}^{+}$(2) in climate experiment.
@@
Figure 10: The inputs and outputs of a ClimateNet predictor after applying various transformations. The plotted $x$ values are visualizations of TMQ. In the outputs, purple indicates background and yellow represents atmospheric river.

### v2

@@ Abstract
Existing symmetry discovery methods predominantly focus on global transformations across the entire system or space, but they fail to consider the symmetries in local neighborhoods. This may result in the reported symmetry group being a misrepresentation of the true symmetry. In this paper, we formalize the notion of local symmetry as atlas equivariance. Our proposed pipeline, automatic local symmetry discovery (AtlasD), recovers the local symmetries of a function by training local predictor networks and then learning a Lie group basis to which the predictors are equivariant. We demonstrate AtlasD is capable of discovering local symmetry groups with multiple connected components in top-quark tagging and partial differential equation experiments. The discovered local symmetry is shown to be a useful inductive bias that improves the performance of downstream tasks in climate segmentation and vision tasks. Our code is publicly available at https://github.com/Rose-STL-Lab/AtlasD.
@@ 2 Related Work
Equivariant Neural Networks. Equivariant neural networks use known symmetry as an inductive bias when fitting a model. Group equivariant CNN extends the translational equivariance of a CNN to rotations and reflection using group theory (Cohen & Welling, 2016). Other works focus on designing networks that are equivariant to a wider range of transformations; particularly, $\mathrm{E}(2)$ transformations on the Euclidean plane (Weiler & Cesa, 2019), rotations on the sphere (Cohen et al., 2018), Lorentz group transformations (Gong et al., 2022), and $\mathrm{E}(n)$ transformations in higher-dimensional spaces (Satorras et al., 2021). These works focus on global transformations and require prior knowledge of the symmetry. In contrast, our pipeline focuses on discovering local symmetry.
@@
To deal with situations where symmetries are only partially present in systems, some works introduce the notion of approximately equivariant neural networks. Implementing such networks can be done in various manners, including modification of convolutional layers (Wang et al., 2022; van der Ouderaa et al., 2022) and addition of non-equivariant residual terms (Finzi et al., 2021). Other works seek to match model and data equivariance error more closely by learning the degree of equivariance from the data (Veefkind & Cesa, 2024a; Romero & Lohit, 2022a; Veefkind & Cesa, 2024b; van der Ouderaa et al., 2023). Petrache & Trivedi (2023) gives results about the generalizability of approximately equivariant networks, whereas Wang et al. (2023) provides a theoretical understanding of equivariant networks in systems with varying degrees of symmetry. While approximate equivariance is related to local symmetry in that they both deal with situations where the global symmetry may not be fully present, local symmetry is a more generalized notion as it can be applied to arbitrary manifolds.
@@
Symmetry discovery works also consider slightly varied problem settings. Some authors seek to find a subset of possible symmetries (Benton et al., 2020; Romero & Lohit, 2022b). Others consider the case where the group acts on the latent space instead of the feature space (Yang et al., 2024; Gabel et al., 2023; Keurti et al., 2023; Koyama et al., 2024). Rather than focusing solely on the symmetry group, van der Ouderaa et al. (2024); Hou et al. (2024) learn the conserved quantities of systems. Still, these papers mainly focus on global transformations. More relevant to local symmetry discovery is the work by Decelle et al. (2019), which attempts to see if two datapoints are related by a particular local transformation. This is distinct from AtlasD, where we characterize the local symmetry group in an interpretable manner.
Comparison with State-of-the-Art. LieGAN, LieGG, and AtlasD are all capable of discovering continuous equivariances, but LieGG requires significant memory and LieGAN uses adversarial methods, which can be complex to train. Moreover, while LieGG cannot discover any discrete symmetries and LieGAN can only discover those with positive determinants, AtlasD proves capable of discovering orientation-reversing discrete symmetries. The largest difference, however, is the implicit hypothesis space of each model. LieGG and LieGAN only consider global symmetries, but AtlasD can also discover local ones. The hypothesis space is a crucial component of a discovery method as it determines the set of symmetries that can be recovered.
@@ 4 AtlasD: Automatic Local Symmetry Discovery
Existing discovery methods are fundamentally limited in that they ignore local symmetries. To address this problem, we first formulate atlas equivariance as a definition of local symmetry. Then, we detail our methodology for discovering local symmetry in the form of a Lie algebra basis and component group. Finally, we highlight theoretical connections to existing work as well as implementation notes.
@@ 4.1 Atlas Equivariance
To provide an intuition of local symmetry, we highlight the heat equation on a torus in Figure 2. Consider the time-stepping function that evolves the current state of the system for some fixed time interval. If we focus only on a neighborhood of the input feature field and its corresponding neighborhood in the output field, a local rotation in the input results in an identical local rotation in the output.
To define local symmetry formally, assume we have a function map $\Phi$ that transforms an input feature field $F_{\text{in}}:{\mathcal{M}}\to\mathbb{R}^{\text{d}_{\text{in}}}$ to an output field $F_{\text{out}}:{\mathcal{M}}\to\mathbb{R}^{\text{d}_{\text{out}}}$. Then, suppose $\mathcal{A}$ is an atlas on ${\mathcal{M}}$ given by a finite collection of charts $\{(U_{c},\varphi_{c})\}_{c=1}^{N}$. For any feature field $F$ on ${\mathcal{M}}$, we can relocate the subset of $F$ contained in the neighborhood $U_{c}$ to a flat Euclidean space by pulling back over $\varphi_{c}^{-1}$:
@@
We define local (atlas) equivariance for functions $\Phi$ where the output signal depends locally on the input signal. We formalize this as $\mathcal{A}$ locality. The intuitive notion of a local function is that, under an appropriate choice of atlas, we can fully reconstruct the output field in any individual chart $U_{c}$ solely from the input field along the same chart. Thus, we say $\Phi$ is $\mathcal{A}$ atlas local if we are able to decompose it into various $\Phi_{c}$, where $\Phi_{c}$ is a map between the pullback of the $c$th chart of the input and output feature fields. We sometimes refer to $\{\Phi_{c}\}$ as the localized functions of $\Phi$. Formally,
@@
Here, the restriction of the feature field to the subset $\varphi_{c}(U_{c})$ indicates that we are not particular about the output of $\Phi_{c}$ outside the projected chart. We also note that atlas locality is related to the idea of sheaf morphisms. In Appendix A, we show that the former is a weaker condition than the latter.
For these atlas local functions, it is possible to consider the symmetry transformations that operate within local neighborhoods. We formalize this notion of local symmetry as follows.
@@ 4.2 Atlas Equivariance Discovery
In our problem setup, given an unknown task function $\Phi:\mathcal{X}\to\mathcal{Y}$ between feature fields on the same manifold, we have a dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$. We assume that a suitable atlas $\mathcal{A}$ for the problem is known. We further assume that the dataset is large enough such that ground truth symmetry is represented in any chart (i.e., all members of any orbit are present). We aim to find the maximal matrix Lie group to which $\Phi$ is $\mathcal{A}$ atlas equivariant.
We propose AtlasD, automatic local symmetry discovery, to tackle the problem. First, we must approximate the individual localized functions $\Phi_{c}$ with neural networks or other differentiable oracles. The exact method is unique to each task, but is generally a simple regression problem. Further details are available in Appendix C. In contrast to global discovery techniques, we emphasize that the individual neural networks are localized maps rather than functions over the entire manifold.
Next, we find the symmetry group of the localized functions. Note that there are important differences between our procedure and global symmetry discovery. In our setting, a group element must only act on a local chart, rather than the full space. Likewise, a core tenet of local symmetry is that we can apply different transformations to different regions of the feature field. Care must be taken, therefore, that the actions on different charts are truly independent.
@@
  1. Train a predictor network $\Phi_{c}$ for each chart $(U_{c},\varphi_{c})$ under loss $\mathcal{L}\left(\Phi_{c}((\varphi_{c}^{-1})^{*}X_{i}),(\varphi_{c}^{-1})^{*}Y_{i}\right)$
  2. Given the trained predictors $\{\Phi_{c}\}$, discover the Lie algebra basis $\{B_{i}\}$
      Set $g=\exp(\sum_{i=1}^{k}\eta_{i}B_{i})$ where $\eta\sim N_{k}(0,I)$
      Optimize $B$ under loss       $\mathcal{L}(g\cdot\Phi_{c}((\varphi_{c}^{-1})^{*}X_{i}),\Phi_{c}(g\cdot(\varphi_{c}^{-1})^{*}X_{i}))$
  3. Given $\{\Phi_{c}\}$ and $\{B_{i}\}$, find the coset representatives $\{C_{\ell}\}$
      Set $g=\text{normalize}(C_{\ell})$
      Optimize $C_{\ell}$ under loss       $\mathcal{L}(g\cdot\Phi_{c}((\varphi_{c}^{-1})^{*}X_{i}),\Phi_{c}(g\cdot(\varphi_{c}^{-1})^{*}X_{i}))$
      Filter duplicate cosets using $B$
@@ 4.3 Connection to Gauge Equivariant CNN
A related notion of local symmetry is introduced by Cohen et al. (2019) when defining gauge equivariant CNNs. In short, gauge equivariance implies that one should be able to arbitrarily orient the local coordinate systems used to define input features and compute convolutions. Hence, it is a property of the deep learning model, rather than the task function itself. This is a notable difference compared to our work, where the atlas equivariance group is intrinsic to the system and can be discovered.
@@ 5.1 Global Symmetry Comparison
Figure 4: AtlasD discovers $\mathrm{O}^{+}(1,3)$ in the top tagging task. Here, each red and blue heatmap denotes a Lie algebra basis element. For each generator, the value of its entries are depicted by the individual colors. When read in row-major order, generators 0, 1, 2 correspond to $\mathrm{SO}(3)$ rotation and generators 4, 5, 6 indicate boosts. The pink and green heatmap in the bottom-right displays the computed invariant metric.
Figure 5: AtlasD discovers two cosets in the top tagging experiment: a representative from the identity and parity component. Each yellow and purple heatmap depicts a coset’s representative in matrix form, where the colors denote the values of that matrix’s entries.
@@
First, we fit a predictor network to the dataset using a $3$-layer MLP. Recall that at this point, we have no information about symmetries of the system and thus this is a non-equivariant model. The trained predictor is assumed to be accurate enough such that its symmetries agree with those of the dataset. Hence, we now seek to find the symmetries of the predictor.
The goal is to classify between top quark and lighter quarks jets present in the Top Quark Tagging Reference Dataset (Kasieczka et al., 2019). The dataset contains 2M observations, consisting of four-momentum of up to $200$ particle jets. The classification is invariant to the entire Lorentz group $\mathrm{O}(1,3)$, which we will try to discover.
We use our infinitesimal generator discovery pipeline to learn the invariances of the predictor. We seed our basis with $7$ generators. In Figure 4, we show that the discovered basis matches closely with that of $\mathrm{SO}^{+}(1,3)$, the identity component of the Lorentz group. Moreover, computing the invariant tensor using the method from Yang et al. (2023), we find that the invariant tensor has a cosine correlation of $0.9996$ with the ground truth Minkowski tensor $\text{diag}(-1,1,1,1)$. This is a strong result that is slightly superior to LieGAN’s cosine correlation of $0.9975$.
We then try to discover the various cosets of the symmetry group. We seed our discovery process with $K=64$ matrices. In the dataset, the time component of all momenta are positive, and hence it is difficult to find the time-reversal generator of the Lorentz group. In the last two heatmaps of Figure 4, we discover a parity transformation, which means that the entire learned symmetry group is $\mathrm{O}^{+}(1,3)$. In contrast, LieGAN cannot discover orientation-reversing transformations and hence only reports the identity component $\mathrm{SO}^{+}(1,3)$.
@@ 5.2 Partial Differential Equation
(a) Example input and output field
(b) Charts in each atlas
Figure 6: Illustration of PDE experiment. The top row depicts an example input and output scalar field for the heat experiment. The bottom row shows the two atlases used for the experiment. Each blue or red square represents an individual chart.
@@
Figure 7: PDE discovered symmetry. The results for the first and second atlas are depicted in the top and bottom rows, respectively. In each case, the leftmost entry is the discovered infinitesimal generator, and the right two columns are the discovered cosets.
We also run the global symmetry discovery baseline LieGG in a similar setup. The full details are deferred to Section D.4. In short, LieGG fails to discover any global symmetry due to the heat source (Figure 8).
Figure 8: Singular values of generators discovered by LieGG in modified PDE experiment. We run twice: once without the rectangular heat source (blue) and once with the heat source (red). The run without the heat source provides a reference for singular values in the case that there is global symmetry.
@@ 5.3 MNIST on Sphere
Figure 9: MNIST experiment setup. (a) The input feature fields (top) are given by three digits rotated and then projected onto the equator of a sphere. To construct the output feature fields (bottom), the model must label each pixel as either background or its numeric value if it is a part of a number. (b) We highlight two of the charts used in our atlas.
To highlight the benefits of using our learned results in downstream models, we design a projected MNIST segmentation task. In this experiment, we project three digits from the MNIST dataset onto a sphere (Figure 9). Before projection, each image is randomly rotated up to $60$ degrees clockwise or counterclockwise. The goal of the model is to classify each pixel as either the background or its numeric value. Although the rotation of the digits adds a local symmetry, there is no continuous global symmetry since the position of each of the three digits is fixed. For this problem, we construct an atlas by assigning a single chart to the region of each digit. This is an admittedly idealized setup, but our main purpose is to demonstrate the full pipeline.
We then train a predictor for each of the three charts using CNNs. In the discovery process, we seed our model with a single infinitesimal generator. To demonstrate the benefit of considering local symmetry, we compare our results against a modified LieGAN that represents global symmetries as subgroups of $\mathrm{SO}(3)$.
@@ 5.4 ClimateNet
Figure 11: AtlasD discovers $\mathrm{GL}^{+}$(2) in climate experiment.
@@
We use an atlas that has $4$ charts spread through the surface of the Earth. When we seeded our model with $1,2,$ or $3$ infinitesimal generators, we find that the resultant basis is not similar across consecutive runs. This suggests that the symmetry group is actually $4$-dimensional. To confirm this, we plot the output of a chart predictor $f$ after applying various linear actions in Figure 12. The figure highlights that the predictor is mildly equivariant to a wide range of actions. In fact, Figure 11 demonstrates comparable magnitudes of the $4$ generators. All of these suggest that the atlas equivariance group is $\mathrm{GL}^{+}(2)$.
@@
Figure 12: The inputs and outputs of a ClimateNet predictor $f$ after applying various transformations. The plotted $x$ values are visualizations of TMQ. In the outputs, purple indicates the background, and yellow represents the atmospheric river.